Simultanius ordinary diff. Eqn dx/x^2-y^2-z^2=dy/2xy=dz/2zx
1
Expert's answer
2020-12-21T19:21:56-0500

"\frac{dx}{x^2-y^2-z^2}=\frac{dy}{2xy}=\frac{dz}{2zx}"


"\frac{dy}{y}=\frac{dz}{z}"

"lny=ln(c_1z)"

"\frac{y}{z}=c_1"


"\frac{dx}{x^2-(c_1z)^2-z^2}=\frac{dz}{2zx}"

"\frac{2zx}{x^2-(c_1z)^2-z^2}=\frac{dz}{dx}"

This is a homogeneous DE.

"z=tx, z'=t'x+t"

"t'x+t=\frac{2tx^2}{x^2-t^2x^2(c_1+1)}=\frac{2t}{1-t^2(c_1+1)}"

"t'x=\frac{2t}{1-t^2(c_1+1)}-t=\frac{t+t^3(c_1+1)}{1-t^2(c_1+1)}"


"\intop\frac{1-t^2(c_1+1)}{t+t^3(c_1+1)}dt=\intop\frac{dx}{x}"


"\frac{1-t^2(c_1+1)}{t+t^3(c_1+1)}=\frac{1-t^2(c_1+1)}{t(1+t(c_1+1)^{1/3})(1-t(c_1+1)^{1/3}+t^2(c_1+1)^{2/3})}"


"\frac{1-t^2(c_1+1)}{t(1+t(c_1+1)^{1/3})(1-t(c_1+1)^{1/3}+t^2(c_1+1)^{2/3})}=\frac{A}{t}+\frac{B}{1+t(c_1+1)^{1/3}}+\frac{Ct+D}{1-t(c_1+1)^{1/3}+t^2(c_1+1)^{2/3}}"


"A(1+t^3(c_1+1))+Bt(1-t(c_1+1)^{1\/3}+t^2(c_1+1)^{2\/3})+(Ct+D)t(1+t(c_1++1)^{1\/3})=1-t^2(c_1+1)"


"A+At^3(c_1+1)+Bt-Bt^2(c_1+1)^{1\/3}+Bt^3(c_1+1)^{2\/3}+Ct^2+Ct^3(c_1+1)^{1\/3}+"

"+Dt+Dt^2(c_1+1)^{1\/3}=1-t^2(c_1+1)"


"A=1"

"B+D=0"


Let "(c_1+1=k^3)" , then:


"-Bk+C+Dk=-k^3"

"C-2Bk=-k^3"

"Ak^3+Bk^2+Ck=0"

"C=-k^2-Bk"

"-k^2-Bk-2Bk=-k^3"

"B=\frac{k^2-k}{3}, D=\frac{k-k^2}{3}"

"C=\frac{-2k^2-k^3}{3}"


"\intop\frac{1-t^2(c_1+1)}{t+t^3(c_1+1)}dt=\intop(\frac{1}{t}+\frac{k^2-k}{3(1+kt)}+\frac{k-k^3-t(2k^2+k^3)}{3(1-kt+k^2t^2)})dt"


"\intop\frac{k-k^3-t(2k^2+k^3)}{3(1-kt+k^2t^2)}dt=\frac{k}{3}\intop\frac{1-k^2-kt(k+2)}{1-kt+k^2t^2}dt="


"=\frac{k^2-1}{3\sqrt{3}}arctan{\frac{2kt}{\sqrt{3}}}-\frac{1}{6\sqrt{3}}ln(t^2-t/k+1/k^2)+\frac{1}{3k\sqrt{3}}arctan{\frac{2kt}{\sqrt{3}}}"


"lnt+\frac{k-1}{3}ln(1+kt)+\frac{k^2-1}{3\sqrt{3}}arctan{\frac{2kt}{\sqrt{3}}}-\frac{1}{6\sqrt{3}}ln(t^2-t/k+1/k^2)+\frac{1}{3k\sqrt{3}}arctan{\frac{2kt}{\sqrt{3}}}="

"=lnx+c_2"


"ln(z/x)+\frac{k-1}{3}ln(1+kz/x)+\frac{k^2-1}{3\sqrt{3}}arctan{\frac{2kz}{x\sqrt{3}}}-\frac{1}{6\sqrt{3}}ln((z/x)^2-z/(kx)+1/k^2)++\frac{1}{3k\sqrt{3}}arctan{\frac{2kz}{x\sqrt{3}}}-lnx=c_2"